The perimeter of a rectangular goat pen is 26 meters. It is 5 meters wide. How long is it?

Answer by macston(5194)

P=2(L+W)
26 meters=2(L+5 meters) divide both by 2
13 meters=L+5 meters subtract 5 from each
13 meters-5 meters=L
8 meters=L ANSWER Length=8
CHECK
P=2(L+W)
26 meters=2(8 meters+5 meters)
26 meters=2(13 meters)
26 meters=26 meters